# L2R: Learning to Reduce Search Space for Generalizable Neural Routing Solver

### v1

@@ 1 Introduction
The Vehicle Routing Problem (VRP) is one of the core problems in Operations Research with significant practical implications in domains such as logistics, supply chain management, and express delivery (Tiwari & Sharma, 2023; Sar & Ghadimi, 2023). Efficient routing optimization is crucial for enhancing delivery performance and reducing operational costs. Traditional heuristic algorithms, such as LKH3 (Helsgaun, 2017) and HGS (Vidal, 2022), have demonstrated strong capabilities in solving VRPs with diverse constraints. However, these methods face two fundamental limitations: (1) their design requires extensive domain expertise to craft problem-specific rules, and (2) their computational complexity scales poorly with instance size due to the NP-hard nature of VRPs. These challenges are particularly acute for large-scale instances (e.g., more than 10,000 nodes), where existing algorithms often fail to provide practical solutions with a reasonable runtime.
@@
In recent years, neural combinatorial optimization (NCO) has emerged as a promising paradigm for solving complex problems like VRPs, which eliminates the need for time-consuming and handcrafted algorithm design by experts (Bengio et al., 2021; Li et al., 2022). These methods automatically learn problem-specific patterns through training frameworks such as supervised learning (SL) (Vinyals et al., 2015; Luo et al., 2023; Drakulic et al., 2023) or reinforcement learning (RL) (Bello et al., 2016; Kool et al., 2019; Zhou et al., 2024). A well-trained NCO model can directly construct approximate solutions without explicit search, offering a promising direction for real-time VRP solving. However, SL-based methods face a critical limitation due to the difficulty of obtaining high-quality labeled data (e.g., nearly optimal solutions) for large-scale NP-hard problems. In contrast, RL-based methods do not require labeled data and have demonstrated strong performance on small-scale instances (e.g., 100 nodes) (Kwon et al., 2020; Xin et al., 2021; Kim et al., 2022). Nevertheless, their effectiveness diminishes significantly on large-scale instances (e.g., 10,000 nodes), primarily due to the exponentially growing search space and the challenge of sparse reward.
To address the scalability challenges, search space reduction (SSR) has gained increasing attention as a scalable strategy. As listed in Table 1, existing SSR techniques can be broadly categorized into two types: static and dynamic. Static SSR performs a one-time pruning at the beginning of the optimization process, offering computational efficiency. However, it often requires additional search procedures (e.g., Monte Carlo Tree Search for TSP) to achieve high-quality solutions (Fu et al., 2021; Qiu et al., 2022; Sun & Yang, 2023). In contrast, dynamic SSR (Fang et al., 2024; Gao et al., 2024; Wang et al., 2024) adaptively adjusts the candidate node set at each construction step based on real-time problem states, enabling more effective search space reduction for constructive NCO methods. Despite their advantages, existing dynamic SSR methods are fundamentally constrained by their reliance on distance-based node selection, which struggle to generalize to large-scale instances, particularly those with non-uniform node distributions.
  • We comprehensively analyze the key limitation of existing distance-based SSR methods and then propose a novel learning-based framework with hierarchical static and dynamic SSR for solving large-scale VRPs.
  • In particular, we develop an RL-based approach to reduce search space and select candidate nodes at each construction step, significantly reducing computational overhead without compromising solution quality.
@@
Neural Routing Solver Static Dynamic Training Generalizable
(Ordered by year of publication) SSR SSR Scale Scale
MLPR (Sun et al., 2021) $\checkmark$ $\times$ $100$ 2K
Att-GCN+MCTS (Fu et al., 2021) $\checkmark$ $\times$ $50$ 10K
DIMES (Qiu et al., 2022) $\checkmark$ $\times$ 10K 10K
DIFUSCO (Sun & Yang, 2023) $\checkmark$ $\times$ 10K 10K
T2T (Li et al., 2023) $\checkmark$ $\times$ 1K 1K
BQ (Drakulic et al., 2023$\ddagger$ $\times$ Distance-based $001$K
ELG (Gao et al., 2024) $\times$ Distance-based $100$ 7K
DAR (Wang et al., 2024) $\times$ Distance-based $500$ 11K
INViT (Fang et al., 2024) $\times$ Distance-based $010$K
@@
    BQ (Drakulic et al., 2023) limits the sub-graph to the $250$ nearest neighbors of the current node when facing large-scale instances.
2 Related Work
2.1 NCO without Search Space Reduction
Most neural combinatorial optimization (NCO) models are trained on small-scale instances (e.g., 100 nodes) without search space reduction and achieve strong performance on instances of similar size. However, their effectiveness significantly diminishes when applied to larger instances (e.g., those exceeding 1,000 nodes) (Kool et al., 2019; Kwon et al., 2020). Some approaches incorporate additional search procedures, such as 2-opt (Deudon et al., 2018) and active search (Bello et al., 2016; Hottung et al., 2022), to address this limitation. While these techniques can improve solution quality, they are still computationally expensive for large-scale instances. Another line of research focuses on training NCO models directly on larger-scale instances (e.g., up to 500 nodes) to enhance generalization (Cao et al., 2021; Zhou et al., 2023, 2024). However, this approach incurs prohibitive computational costs due to the exponentially growing search space. Alternatively, some methods simplify large-scale VRPs through decomposition policies (Kim et al., 2021; Li et al., 2021; Hou et al., 2022; Pan et al., 2023; Zheng et al., 2024). Although effective, these methods often overlook the dependency between decomposition policies and subsequent solvers, resulting in suboptimal solutions. Moreover, the reliance on expert-designed policies limits their practicality for real-world applications.
2.2 NCO with Static Search Space Reduction
To address the scalability challenges of NCO methods, static search space reduction (SSR) has been proposed as a computationally efficient approach. These methods perform one-time pruning at the beginning of the optimization process, significantly reducing the problem size. For example, Sun et al. (2021) develop a static problem reduction technique to eliminate unpromising edges in large-scale TSP instances. Recently, heatmap-based approaches have gained popularity for solving large-scale TSPs, where models are trained to predict the probability of each edge belonging to the optimal solution. To handle large-scale instances, these methods often incorporate graph sparsification (Qiu et al., 2022; Sun & Yang, 2023; Li et al., 2023) or pruning strategies (Fu et al., 2021; Min et al., 2023) to reduce the search space. While static SSR is computationally efficient, it typically requires additional well-designed search procedures (e.g., Monte Carlo Tree Search for TSP) to achieve high-quality solutions, which might be more important for the optimization process (Xia et al., 2024).
2.3 NCO with Dynamic Search Space Reduction
Dynamic search space reduction (SSR) has also been explored as a promising approach to address the scalability challenges of NCO methods. These methods adaptively prune the search space to a small set of candidate nodes at each construction step, typically based on the distance to the last visited node. The final node selection can be guided by either the original policy augmented with auxiliary distance information (Wang et al., 2024) or a well-designed local policy (Gao et al., 2024). Additionally, recent work by Fang et al. (2024) and Drakulic et al. (2023) directly selects the next node from the candidate set using NCO models. While dynamic SSR can more efficiently reduce the search space for constructive NCO, its reliance on distance-based node selection is inappropriate for large-scale instances, especially those from non-uniform distributions.
@@ 2.3 NCO with Dynamic Search Space Reduction
Figure 2: Impact of distance-based search space reduction on solution optimality. (a)-(d) Optimal solutions for TSP1000 instances under four distribution patterns: uniform, cluster, explosion, and implosion. Restricting the search space to the k-nearest neighbors (k = 20) will lead to suboptimal routes.
Figure 3: The effects of distance-based search space reduction on optimality ratio (TSP100).
@@ 3 Shortcomings of Distance-based Search Space Reduction
Constructive NCO methods solve routing problems by iteratively selecting nodes to build a solution. In theory, each step of this process should consider all unvisited nodes to guarantee optimality. However, for large-scale instances, this exponentially growing search space could quickly become computationally intractable. While current distance-based SSR methods can significantly improve computational efficiency, they also introduce critical limitations to NCO. In this section, we systematically analyze their shortcomings through two key perspectives: (1) the degradation of solution optimality and (2) the impact for the NCO solver.
Degradation of Solution Optimality
@@
The fundamental limitation of distance-based search space reduction lies in its tendency to prune globally optimal nodes during sequential solution construction. As demonstrated in Figure 2, restricting candidate nodes to the $k$-nearest neighbors forces the algorithm to ignore critical long-range connections necessary for optimal routing. This over-pruning effect accumulates systematically across construction steps, ultimately compromising solution quality. The degradation is particularly severe in instances from non-uniform distributions, where optimal paths inherently rely on non-local node selections.
To quantify the impact of search space reduction size $k$ on solution optimality, we analyze the probability of retaining optimal nodes when restricting candidate selections to the $k$-nearest neighbors. For each of the four distributions (uniform, clustered, explosion, implosion), we generate 2,000 TSP100 instances and compute their optimal solutions using LKH3 (Helsgaun, 2017). We then measure the optimality ratio, defined as the proportion of construction steps where the optimal node resides within the $k$-nearest candidates.
As shown in Figure 3, the optimality ratio monotonically increases with $k$, and each distribution exhibits a critical threshold $k^{*}$ such that all optimal nodes are preserved when $k\geq k^{*}$. This implies that all nodes outside the $k^{*}$-nearest neighbors are redundant for optimal routing. However, when $k$ is small, the optimality ratio drops significantly. Moreover, a large $k^{*}$ indicates that search space reduction cannot significantly reduce the computational overhead, thereby failing to address the scalability challenge of solving large-scale routing problems.
Figure 4: The effect of distance-based search space reduction on the gap in LEHD model (Luo et al., 2023). Here $\Delta Gap=Gap_{\text{reduced}}-Gap_{\text{original}}$.
Impact for Constructive NCO
To evaluate the impact of search space reduction on NCO performance, we use the well-known LEHD method (Luo et al., 2023) as a representative example, restricting the search space to the $k$ nearest nodes from the last visited node at each construction step. We quantify the impact by measuring the performance gap $\Delta Gap$, defined as the difference between the optimality gap obtained by LEHD with a reduced search space ($Gap_{\text{reduced}}$) and the original LEHD without search space reduction ($Gap_{\text{original}}$).
The results in Figure 4 illustrate the performance on TSP instances of varying sizes under different search space reduction levels ($k$). A key observation is the existence of a critical threshold $k^{*}$ for each problem size, beyond which LEHD with $k\geq k^{*}$ achieves a $0\%$ performance gap with the original LEHD, indicating no loss in solution quality with SSR. However, these critical thresholds vary significantly across problem sizes, and using a small $k<k^{*}$ leads to a substantial performance degradation. These findings highlight the potential of search space reduction while underscoring the need for advanced reduction methods capable of efficiently handling large-scale problems with a small $k$.
@@ 4.1 Static Reduction
The static reduction stage initiates our hierarchical framework by pruning the original fully connected graph $G=(V,E)$ into a sparse topology $G^{\prime}=(V,E^{\prime})$. For each node $v_{i}\in V$, we compute pairwise Euclidean distances $\{d_{ij}\}_{j=1}^{N}$ and eliminate connections to nodes in the farthest $\alpha$-percentile. Formally, the sparse edge set is defined as:
$E^{\prime}=\bigcup_{v_{i}\in V}\left\{e_{ij}\mid\mathrm{rank}(d_{ij})\leq(1-\alpha)|V|\right\},\quad\alpha\in[0,1]$ (1)
where $\mathrm{rank}(d_{ij})$ denotes the ascending order of distances from $v_{i}$ (i.e., $\mathrm{rank}(d_{ij})=1$ for the closest neighbor). Through empirical analysis across diverse scales and node distributions, we use $\alpha=10\%$ as the threshold in this work, which can efficiently reduce the computational overhead without compromising solution optimality. Crucially, this stage operates as a one-time preprocessing step, incurring no runtime overhead during subsequent solving phases.
@@ 4.2 Learning-based Reduction
To address this critical issue, we develop a learning-based model to dynamically evaluate the potential of feasible nodes and adaptively reduce the search space at each construction step. Our proposed model ensures efficiency through a lightweight structure containing only an embedding layer and an attention layer. Their implementations are detailed in the following.
@@
Given an instance $S=\{\mathbf{s}_{i}\}_{i=1}^{N}$ with node features $\mathbf{s}_{i}\in\mathbb{R}^{d_{x}}$ (e.g., city coordinates in TSP), we first project these features into $d$-dimensional embeddings through a shared linear transformation for each node:
@@
At $t$-th step with partial solution $(\pi_{1},\dots,\pi_{\pi_{t-1}})$, we adopt the setting from Drakulic et al. (2023) to represent the current partial solution using the initial node embedding $\mathbf{h}_{\pi_{1}}^{D}$ and latest node embedding $\mathbf{h}_{\pi_{t-1}}^{D}$. In addition, following Kool et al. (2019) and Kwon et al. (2020), we define a context embedding of the current partial solution:
$\mathbf{h}_{(C_{D})}^{t}=W_{\text{first}}\mathbf{h}_{\pi_{1}}^{D}+W_{\text{last}}\mathbf{h}_{\pi_{t-1}}^{D},$ (3)
@@
where $W^{K},W^{V}\in\mathbb{R}^{d\times d}$ are learnable projection matrices. The context embedding then interacts with these projections through the attention operator:
@@
Finally, the potential scores $\{o_{i}|i\in\mathcal{A}_{t}\}\in\mathbb{R}^{1\times|\mathcal{A}_{t}|}$ can be computed as
$o_{i}=\sigma\left(\frac{\hat{\mathbf{h}}_{(C_{D})}^{t}(W^{C}\mathbf{h}_{i}^{D})^{\mathrm{T}}}{\sqrt{d_{k}}}\right)-d_{t-1,i},\ \forall i\in\mathcal{A}_{t},$ (6)
where $\sigma$ represents the Sigmoid function, $W^{C}\in\mathbb{R}^{d\times d}$ is a learnable parameter matrix, $d_{k}$ is the dimension for matrix $K_{\mathcal{A}_{t}}^{D}$, and $d_{t-1,i}$ denotes the distance between the current node $\pi_{t-1}$ and each feasible node $i$ with $i\in\mathcal{A}_{t}$, which is normalized into $[0,1]$ via division by $\sqrt{2}$ (i.e., the diagonal distance of unit square). Each score $o_{i}$ quantifies the potential of node $i\in\mathcal{A}_{t}$. Then we retain the top-$k$ nodes with highest potential scores as the candidate nodes and denote the set of all candidate nodes as $\mathcal{N}_{t}$. In this way, we have conducted a learning-based dynamic search space reduction at each construction step. More details can be found in Appendix A.

### v2

# Learning to Reduce Search Space for Generalizable Neural Routing Solver

@@ 1 Introduction
The Vehicle Routing Problem (VRP) is one of the core problems in Operations Research with significant practical implications in domains such as logistics, supply chain management, and express delivery [1, 2]. Efficient routing optimization is crucial for enhancing delivery performance and reducing operational costs. Traditional heuristic algorithms, such as LKH3 [3] and HGS [4], have demonstrated strong capabilities in solving VRPs with diverse constraints. However, these methods face two fundamental limitations: (1) their design requires extensive domain expertise to craft problem-specific rules, and (2) their computational complexity scales poorly with instance size due to the NP-hard nature of VRPs. These challenges are particularly acute for large-scale instances (e.g., more than 10,000 nodes), where existing algorithms often fail to provide practical solutions with a reasonable runtime.
@@
In recent years, neural combinatorial optimization (NCO) has emerged as a promising paradigm for solving complex problems like VRPs, which eliminates the need for time-consuming and handcrafted algorithm design by experts [5, 6, 7]. These methods automatically learn problem-specific patterns through training frameworks such as supervised learning (SL) [8, 9, 10, 11, 12, 13] or reinforcement learning (RL) [14, 15, 16, 17]. A well-trained NCO model can directly construct approximate solutions without explicit search, offering a promising direction for real-time VRP solving. However, SL-based methods face a critical limitation due to the difficulty of obtaining high-quality labeled data (e.g., nearly optimal solutions) for large-scale NP-hard problems. In contrast, RL-based methods do not require labeled data and have demonstrated strong performance on small-scale instances (e.g., 100 nodes) [18, 19, 20]. Nevertheless, their effectiveness diminishes significantly on large-scale instances (e.g., 10,000 nodes), primarily due to the exponentially growing search space and the challenge of sparse reward.
@@
Neural Routing Static Dynamic Training Generalizable
Solver SSR SSR Scale Scale
MLPR [21] $\checkmark$ $\times$ $100$ 2K
Att-GCN+MCTS [22] $\checkmark$ $\times$ $50$ 10K
DIMES [23] $\checkmark$ $\times$ 10K 10K
DIFUSCO [24] $\checkmark$ $\times$ 10K 10K
T2T [25] $\checkmark$ $\times$ 1K 1K
BQ [10$\ddagger$ $\times$ Distance-based $001$K
ELG [26] $\times$ Distance-based $100$ 7K
DAR [27] $\times$ Distance-based $500$ 11K
INViT [28] $\times$ Distance-based $010$K
@@
    BQ [10] limits the sub-graph to the $250$ nearest neighbors of the current node when facing large-scale instances.
To address the scalability challenges, search space reduction (SSR) has gained increasing attention as a scalable strategy. As listed in Table 1, existing SSR techniques can be broadly categorized into two types: static and dynamic. Static SSR performs a one-time pruning at the beginning of the optimization process, offering computational efficiency. However, it often requires additional search procedures (e.g., Monte Carlo Tree Search for TSP) to achieve high-quality solutions [22, 23, 24]. In contrast, dynamic SSR [28, 26, 27] adaptively adjusts the candidate node set at each construction step based on real-time problem states, enabling more effective search space reduction for constructive NCO methods. Despite their advantages, existing dynamic SSR methods are fundamentally constrained by their reliance on distance-based node selection, which struggle to generalize to large-scale instances, particularly those with non-uniform node distributions. We provide a comprehensive literature review of search space reduction approaches in Appendix B.
  • We comprehensively analyze the key limitations of existing distance-based SSR methods and then propose a novel learning-based framework with hierarchical static and dynamic SSR for solving large-scale VRPs.
  • In particular, we develop an RL-based approach to reduce the search space and select candidate nodes at each construction step, significantly reducing computational overhead without compromising solution quality.
@@
Figure 2: Impacts of distance-based search space reduction on solution optimality. (a)-(d) Optimal solutions for TSP1K instances under four distribution patterns: uniform, cluster, explosion, and implosion. Restricting the search space to the k-nearest neighbors ($k=20$) will lead to suboptimal routes. The red, green, and blue nodes indicate the current node, the next visiting node, and $k$ nearest neighbors, respectively.
@@ 2 Shortcomings of Distance-based Search Space Reduction
The fundamental limitation of distance-based search space reduction lies in its tendency to prune globally optimal nodes during solution generation. As demonstrated in Figure 2, restricting candidate nodes to the $k$-nearest neighbors forces the algorithm to ignore critical long-range connections necessary for optimal routing. This over-pruning effect accumulates systematically throughout the solving process, ultimately compromising solution quality. In this section, we systematically analyze these shortcomings through the impact on the optimality gap of the classic solver and the impact on constructive NCO methods across diverse architectures.
Impact on Optimality Gap
Table 2: Optimality gap comparison of LKH-3 on unreduced and reduced TSPLIB instances.
Instance Scale w/o D-SSR w/ $k=10$ w/ $k=20$
dsj1000 1,000 0.00% 7.27% 6.85%
pr1002 1,002 0.00% 3.14% 1.06%
d1291 1,291 0.00% 26.71% 9.59%
fl1400 1,400 0.02% 64.72% 54.29%
fl1577 1,577 0.01% 48.14% 38.19%
d1655 1,655 0.00% 20.41% 8.74%
rl1889 1,889 0.00% 27.94% 8.45%
d2103 2,103 0.01% 17.62% 6.10%
fl3795 3795 0.06% 56.87% 53.38%
rl5915 5,915 0.03% 17.99% 2.58%
rl5934 5,934 0.03% 27.24% 6.40%
rl11849 11,849 0.00% 11.01% 1.90%
usa13509 13,509 0.01% 20.06% 5.87%
Avg.gap 0.01% 26.86% 15.64%
To evaluate the impact of distance-based search space reduction on final optimization performance, we adopt the widely-used LKH-3 algorithm [3] as our benchmark solver and conduct experiments on the TSPLIB dataset [29]. Given that LKH-3 is not a constructive heuristic, we restrict the search space for each instance by pruning the original fully connected graph into a sparse topology. Specifically, only connections to the $k$ nearest nodes are retained for each node in our experiments.
The results in Table 2 illustrate the performance across TSPLIB instances of varying scales under different levels of search space reduction ($k$). A key observation is that while LKH-3 achieves near-optimal solutions in an unreduced search space, aggressive pruning (i.e., using small $k$ values) substantially degrades performance. This degradation is particularly pronounced in instances where optimal routes inherently depend on non-local node selections, which may be eliminated by such aggressive pruning strategies. We provide detailed solution visualizations comparing three different results in Section C.1.
Impact on Constructive NCOs
For constructive NCO methods, each construction step should consider all feasible nodes to guarantee optimality. However, since VRPs are classic NP-hard problems, the computational complexity grows exponentially with problem size. For large-scale instances, directly obtaining high-quality solutions remains extremely challenging in an unreduced search space. While distance-based SSR methods [10, 28] improve computational efficiency, they also introduce critical limitations in their failure to include a small but essential subset of nodes within their candidate set. The experimental results shown in Section C.2 highlight the potential of search space reduction while underscoring the need for advanced reduction methods capable of efficiently handling large-scale problems with a small $k$.
@@ 3.1 Static Reduction
The static reduction stage initiates our hierarchical framework by pruning the original fully connected graph $G=(\Omega,E)$ into a sparse topology $G^{\prime}=(\Omega,E^{\prime})$. For each node $i\in\Omega$, we compute pairwise Euclidean distances $\{d_{ij}\}_{j=1}^{N}$ and eliminate connections to nodes in the farthest $\gamma$-percentile. Formally, the sparse edge set is defined as
$E^{\prime}=\bigcup_{i\in\Omega}\left\{e_{ij}\mid\mathrm{rank}(d_{ij})\leq(1-\gamma)|\Omega|\right\},\quad\alpha\in[0,1]$ (1)
where $\mathrm{rank}(d_{ij})$ denotes the ascending order of distances from $v_{i}$ (i.e., $\mathrm{rank}(d_{ij})=1$ for the closest neighbor). Through empirical analysis across diverse scales and node distributions, we use $\gamma=10\%$ as the threshold in this work, which can efficiently reduce the computational overhead without compromising solution optimality. Crucially, this stage operates as a one-time preprocessing step, incurring no runtime overhead during subsequent solving phases.
@@ 3.2 Learning-based Reduction
To mitigate the negative impact of reduction on optimality, we develop a learning-based model to dynamically evaluate the potential of feasible nodes and adaptively reduce the search space at each construction step. Our proposed model ensures efficiency through a lightweight structure containing only an embedding layer and an attention layer. Their implementations are detailed in the following.
@@
Given an instance $S=\{\mathbf{s}_{i}\}_{i=1}^{n}$ with $n$ node features $\mathbf{s}_{i}\in\mathbb{R}^{d_{x}}$ (e.g., city coordinates in TSP), we first project these features into $d$-dimensional embeddings through a shared linear transformation for each node:
@@
At $t$-th step with partial solution $(\pi_{1},\dots,\pi_{{t-1}})$, we adopt the setting from Kool et al. [16] to represent the current partial solution using the initial node embedding $\mathbf{h}_{\pi_{1}}^{D}$ and latest node embedding $\mathbf{h}_{\pi_{t-1}}^{D}$. In addition, following Kwon et al. [18], we define a context embedding of the current partial solution:
$\mathbf{h}_{C_{D}}^{t}=W_{\text{first}}\mathbf{h}_{\pi_{1}}^{D}+W_{\text{last}}\mathbf{h}_{\pi_{t-1}}^{D},$ (3)
@@
where $W^{K},W^{V}\in\mathbb{R}^{d\times d}$ are learnable projection matrices. The context embedding $\mathbf{h}_{C_{D}}^{t}$ then interacts with these projections through the attention operator:
@@
Compatibility Calculation
Finally, similar to previous work [16, 18, 17], we can compute the compatibilities $\mathbf{u}^{R}=\{u^{R}_{t,i}|i\in\mathcal{A}_{t}\}$ for selecting each node in $\mathcal{A}_{t}$:
$u^{R}_{t,i}=\begin{cases}\xi\cdot\text{tanh}\left(\frac{\hat{\mathbf{h}}_{C_{D}}^{t}(\mathbf{h}_{i}^{D})^{\mathrm{T}}}{\sqrt{d_{k}}}+a_{t-1,i}^{R}\right)&\text{if}\ i\in\mathcal{A}_{t}\\ -\infty&\text{otherwise}\end{cases},$ (6)
where $\xi$ is the clipping parameter, $d_{k}$ is the dimension for matrix $K_{\mathcal{A}_{t}}^{D}$, and $a_{t-1,i}^{R}$ is the adaptation bias between each node $i\in\mathcal{A}_{t}$ and the current node $\pi_{t-1}$, following Zhou et al. [17] (see Appendix D for more details). Finally, the potential scores $\mathbf{o}=\{o_{i}|i\in\mathcal{A}_{t}\}\in\mathbb{R}^{1\times|\mathcal{A}_{t}|}$ can be computed as:
$\mathbf{o}=\mathrm{softmax}(\mathbf{u}^{R}).$ (7)
The score $o_{i}$ quantifies the potential of node $i\in\mathcal{A}_{t}$. Then, we retain the top-$k$ nodes with the highest potential scores as the candidate nodes and denote the set of all candidate nodes as $\mathcal{N}_{t}$. In this way, we conduct a learning-based dynamic search space reduction at each construction step. More details about search space reduction for CVRPs can be found in Appendix E.